Price Quantity Total Cost

2,000 1 1,600

1,800 2 1,800

1,600 3 2,000

What would be the profit for the monopolist if he choose to produce 2 units of the good?

a)     1,600

b)     1,800

c)     3,600

d)    2,000


1
Expert's answer
2021-01-15T18:19:58-0500

Let's first find the total revenue:


"TR=PQ=1800\cdot2=3600."

The profit for the monopolist can be found by subtracting total cost from its total revenue:


"Profit=TR-TC=3600-1800=1800."

Therefore, the answer is (b) 1800.

Answer:

(b) 1800